three consecutive even integers are such that the sum of the smallest and 3 times the second is 38 more than twice the third. what are the integers?
Answer #1
three consecutive even integers are such that the sum of the smallest and 3 times the second is 38 more than twice the third. what are the integers? ---------------------- 1st: 2x-2 2nd: 2x 3rd: 2x+2 -------------------- Equation: 2x-2 +3(2x)= 2x+2+38 2x-2 + 6x = 2x +40 6x = 42 x = 7 ------------- 1st: 14-2 = 12 2nd: 14 3rd: 14+2 = 16 ======================== Cheers, Stan H. Reply to stanbon@comcast.net
answered by: Seana
Answer #2
Let the smallest integer be I Then the 2nd and 3rd are (I + 2) and (I + 4) Since the sum of the smallest and 3 times the second is 38 more than twice the 3rd, then we get: I + 3(I + 2) = 2(I + 4) + 38 I + 3I + 6 = 2I + 8 + 38 4I – 2I = 46 – 6 2I = 40 I = 20 The integers are: {{{highlight_green(20_22_24)}}} ------------ Check ------------ 20 + 3(22) = 2(24) + 38 20 + 66 = 48 + 38 86 = 86
answered by: Demi